Calculate the molar mass of the compound used in a lab in which the following was given: -62.5 grams of an ionic solid was used to lower the freezing point of water to -1.330C.
-This compound contained 3 ions and was dissolved in 1 kg of water.
1
Expert's answer
2020-12-14T14:40:03-0500

Kf for water is 1.853 "K\cdot{kg}\cdot{mol^{-1}}"

The freezing point depression formula is:

"\Delta{T}=iK_fm"


From this formula we can find the molality (m) - moles of a solute in 1 kg of the solvent; i stands for the number of ions produced by 1 molecule of an ionic compound.


"m=\frac{\Delta{T}}{iK_f}=\frac{0-(-1.33)}{1.853\times3}=0.239mol/kg_{solvent}"


There is exactly 1 kg of water.

Therefore, the molar mass is "\frac{62.5g}{0.239mol}=262g/mol"


Answer: 262 g/mol